OpenStudy (anonymous):

Write the area of the square in simplest polynomial form. A = s2 A square with a side length of 2x + 7. A. 4x2 + 49 B. 4x2 + 14 C. 4x2 + 14x + 49 D. 4x2 + 28x + 49

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Multiply it out with FOIL. (2x+7)(2x+7) 1) F - First (Multiply the first numbers in both parenthesis together.) 2x * 2x = 4x^2 2) O - Outer (Multiply the outer numbers in the parenthesis.) 2x * 7 = 14x 3) I - Inner (Multiply the inner numbers in the parenthesis.) 7 * 2x = 14x 4) L - Last (Multiply the last numbers in the parenthesis.) 7 * 7 = 49 5) Now put all those values together, and add like variables. 4x^2+14x+14x+49 = 4x^2+28x+49 Answer: 4x^2+28x+49

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[(a+b)^{2} = a ^{2} +2ab +b ^{2}\]